# Finding Quasars behind the Galactic Plane. IV. Candidate Selection from Chandra with Random Forest

### v1

@@ Abstract
Quasar samples remain severely incomplete at low Galactic latitudes because of strong extinction and source confusion. We conduct a systematic search for quasars behind the Galactic plane using X-ray sources from the Chandra Source Catalog (CSC 2.1), combined with optical data from Gaia DR3 and mid-infrared data from CatWISE2020. Using spectroscopically confirmed quasars and stars from data sets including DESI, SDSS, and LAMOST, we apply a Random Forest classifier to identify quasar candidates, with stellar contaminants suppressed using Gaia proper-motion constraints. Photometric redshifts are estimated for the candidates using a Random Forest regression model. Applying this framework to previously unclassified CSC sources, we identify 6286 quasar candidates, including 863 Galactic Plane Quasar (GPQ) candidates at $|b|<20^{\circ}$, of which 514 are high-confidence candidates. Relative to the previously known GPQ sample, our selected GPQs reach fainter optical and X-ray fluxes, improving sensitivity to low-flux GPQs. In addition, both the GPQ candidates and known GPQs display harder X-ray spectra than the all-sky quasar sample, consistent with increased absorption through the Galactic plane. Pilot spectroscopy confirms two high-confidence GPQ candidates as quasars at spectroscopic redshifts of $z=1.2582$ and $z=1.1313$, and further spectroscopic follow-up of the GPQ sample is underway. This work substantially improves the census of GPQs and provides a valuable target sample for future spectroscopic follow-up, enabling the use of GPQs to refine the reference frames for astrometry and probe the Milky Way interstellar and circumgalactic media with the absorption features of GPQs.
@@ I INTRODUCTION
Quasars are among the most luminous active galactic nuclei (AGNs), powered by accretion onto supermassive black holes, and are detectable out to cosmological distances. Their high luminosities, compactness, and long-lived activity phases make quasars uniquely valuable probes for a wide range of astrophysical and cosmological questions, including the growth of supermassive black holes and their coevolution with host galaxies (e.g., DiMatteo2005; Kormendy_2013; Wu_2015; Banados_2018; Inayoshi_2020; Fan_2023), and the large-scale structure of the Universe through quasar clustering and baryon acoustic oscillations (e.g., Eisenstein_2005; Dawson2013; Blanton2017). In addition, quasars have negligible parallaxes and proper motions, making them ideal reference objects for defining and maintaining celestial reference frames (CRFs; e.g., Ma2009; Gaia_early_2016; Gaia_2023a). Despite major progress in wide-area surveys, however, the completeness of quasar samples remains poor close to the Galactic plane.
@@
Over the past decades, major spectroscopic surveys have dramatically expanded the known quasar population. The Sloan Digital Sky Survey (SDSS; SDSS; York2000) and its extensions established the first homogeneous, large-area quasar samples (Richards2002; Schneider2010; Paris2017; Lyke2020), while complementary efforts such as the 2dF QSO Redshift Survey and the LAMOST quasar survey extended quasar identification to different depths and sky regions (Croom2004; Croom2009; Ai2016; Jin2023; Lyu2025). More recently, the Dark Energy Spectroscopic Instrument (DESI; DESI_DR1) has delivered the largest spectroscopic quasar samples to date, with more than one million confirmed quasars now available for statistical studies. Despite this rapid progress, the sky distribution of known quasars remains highly non-uniform.
@@
In particular, the census of quasars at low Galactic latitudes remains severely incomplete. Most optical and near-infrared surveys avoid regions close to the Galactic plane, where strong interstellar extinction, high stellar densities, and severe source confusion substantially degrade the effectiveness of traditional color-based selection techniques (Schlegel1998; Ross2012; Richards2009). As a consequence, the surface density of spectroscopically confirmed quasars decreases sharply toward the Galactic plane, producing a persistent gap in all-sky quasar samples (Fu_2021; Huo2025). This incompleteness limits the use of quasars as uniform tracers of large-scale structure and as astrometric reference objects, and motivates dedicated searches for quasars in the low-latitude sky.
@@
Quasars behind the Galactic plane (GPQs; $|b|\lesssim 20^{\circ}$), are nevertheless of considerable scientific importance. As distant extragalactic point sources with negligible parallaxes and proper motions, GPQs provide stable reference objects that are essential for constructing and validating celestial reference frames in the Galactic disk, where background sources are currently sparse (Mignard2018; GaiaCollaboration2018; GaiaCollaboration2023). GPQs also serve as bright background beacons for absorption-line studies of the Milky Way interstellar and circumgalactic media, enabling measurements of gas kinematics, dust extinction, and chemical abundances along heavily obscured sightlines that are under-sampled at high Galactic latitudes (e.g., Savage1996; Richter2006; Fu_2021; Fu_2024).
@@
Identifying GPQs is observationally challenging. The Galactic plane hosts a dense and diverse population of stars, binaries, and other Galactic sources, including many X-ray emitters, and these contaminants can overlap quasars in color–color and color–magnitude spaces. As a result, GPQs have been explored in relatively few dedicated studies. Early efforts by Im2007 uncovered dozens of bright QSOs/AGNs at $|b|<20^{\circ}$ using radio and near-infrared criteria. A major advance was achieved by Fu_2021, who applied transfer learning to optical and infrared data from Pan-STARRS1 (PS1; PS1) DR1 and AllWISE (Wright2010; Cutri2013), producing a catalog of roughly 160,000 GPQ candidates within $|b|<20^{\circ}$. Subsequent follow-up spectroscopy confirmed approximately 200 GPQs with high success rates (Fu_2022). Werk2024 later confirmed 72 UV-bright AGNs at $|b|<30^{\circ}$, with 25 newly recognized objects, and the LAMOST GPQ program has recently reported $\sim 1300$ additional GPQs based on LAMOST DR10 (Huo2025). These studies demonstrate that combining multiwavelength data with modern classification approaches is essential for making progress in the Galactic plane.
@@
Machine-learning methods applied to multi-band photometry and astrometry can exploit high-dimensional information to improve quasar selection in crowded fields (Richards2009; Carrasco2015; Gaia_2023a; Fu_2025). X-ray observations provide a powerful and complementary avenue for identifying quasars with extreme properties (Wang2025) or X-ray sources in the Galactic disk (Bao2025). Among current X-ray facilities, the Chandra X-ray Observatory is uniquely advantageous for low-latitude work: its sub-arcsecond angular resolution substantially mitigates source confusion in dense stellar regions, enabling more reliable associations between X-ray sources and their optical/infrared counterparts. The Chandra Source Catalog (CSC; here we use CSC 2.1) therefore offers an excellent basis for constructing GPQ candidates, provided that Galactic X-ray source contamination can be effectively controlled.
@@ II Data
We used Chandra Source Catalog version 2.1 (CSC 2.1; CSC) Master Sources Table to select GPQ candidates. It includes measured properties for 407,806 unique compact and extended X-ray sources and more than 1.3 million individual detections observed with either ACIS or HRC-I observations released publicly before the end of 2021. Most of these sources remain unidentified. Chandra has two focal plane instruments: Advanced CCD Imaging Spectrometer (ACIS) and High Resolution Camera (HRC). The ACIS instrument observes in broad (b): 0.5–7.0 keV, ultrasoft (u): 0.2–0.5 keV, soft (s): 0.5–1.2 keV, medium (m): 1.2–2.0 keV, and hard(h): 2.0–7.0 keV bands. The HRC instrument observes in 0.1–10 keV energy band and is designated as ‘W’ band. The energy flux in each band is determined using aperture photometry. The source count is derived from an elliptical source region and subtracted by the background count in the surrounding region. To convert the count rate to energy flux, the total count rate is summed up and then scaled by the local ancillary response function.
In this work, we use aperture-corrected average net-flux in b, u, s, m, and h bands, which are named as $b$-csc, $u$-csc, $s$-csc, $m$-csc, $h$-csc, respectively. We applied several constraints to the CSC 2.1 data to ensure quality. First, we selected only point sources with the extent flag set to 0 in the catalog. We then applied additional filters based on the following CSC 2.1 quality flags: pileup_flag, sat_src_flag, conf_flag, streak_src_flag (Table 1). Through this process, a total of 364,555 unique high-quality X-ray point sources were selected from CSC 2.1.
@@
[…]
To improve the purity of the GPQ candidate sample, we then remove potential optically extended sources using the following procedure. We refine the point-source selection by applying the corrected BP/RP photometric excess factor, $C^{*}$, derived from Gaia DR3 (Gaia_cut), which provides an additional constraint on source morphology. The original BP/RP excess factor, $C$, is defined as the ratio of the summed BP and RP integrated fluxes to the $G$-band flux:
@@
Because the BP and RP photometric windows are broader than that of the $G$ band, extended sources tend to exhibit systematically larger excess factors than point sources (e.g., Liu2020). The corrected excess factor $C^{*}$ mitigates the color dependence inherent in $C$, providing a more reliable diagnostic of source extension (Gaia_cut). To compress the dynamic range and suppress the influence of extreme outliers, we impose the criterion $\log(1+C^{*})\leq 0.5$, which yields a cleaner point-source sample (e.g., Fu_2025). Finally 36,152 sources are selected with high-quality measurements from CSC 2.1, CatWISE2020, and Gaia DR3. Hereafter, we refer to this sample as the CSC–WISE–GAIA sample.
@@ II.2 The Training Set
A well-defined and representative training set is critical for the supervised classification of quasars and stars, particularly in the crowded and heterogeneous environments near the Galactic plane. Following the construction of the CSC–WISE–GAIA sample described in Section II.1, we compile spectroscopically confirmed quasar and stellar samples from multiple surveys and catalogs to serve as labeled data for training the RF classifier.
We first identify quasars by cross-matching the CSC–WISE–GAIA sample with spectroscopically confirmed quasars from DESI DR1 (DESI_DR1) and SDSS DR18 (SDSS), adopting a matching radius of $1\arcsec$. To improve completeness, this initial quasar sample is further supplemented with confirmed quasars from the Million Quasars Catalog (Milliquas), the LAMOST spectroscopic surveys (LAMOST_01; LAMOST_02), and the Galactic Plane quasar sample identified from LAMOST DR10 by Huo2025. All supplementary quasar catalogs are cross-matched with the CSC–WISE–GAIA sample using a $3\arcsec$ radius to ensure consistent source associations.
For the stellar class, we similarly cross-match the CSC–WISE–GAIA sample with DESI DR1 and SDSS DR18 to identify spectroscopically classified stars. To better represent the diverse Galactic X-ray source population that may contaminate quasar selection, we further include several specialized stellar catalogs. These include Galactic Wolf–Rayet stars (derHucht2001; derHucht2006) and APOGEE-2 stars from SDSS DR16 (Jonsson2020); young stellar objects (YSOs) in open clusters and molecular clouds (e.g., Ozawa2005; Giardino2007; Delgado2011; Povich2011; Rebull2011; Megeath2012); high-mass X-ray binaries (HMXBs) and related Galactic and nearby extragalactic sources (Liu2006; Walter2015; Mineo2012; Sazonov2017); low-mass X-ray binaries (LMXBs) compiled from the Ritter–Kolb catalog and subsequent updates (Ritter2015; Liu2007; Sazonov2020; Kundu2007; Humphrey2008; Zhang2011); and cataclysmic variables (CVs) (Downes2001; Ritter2015). These catalogs are cross-matched with the CSC–WISE–GAIA sample using a matching radius of $3\arcsec$.
Sources that are not matched to any of the above quasar or stellar catalogs are subsequently queried against the SIMBAD database (Wenger2000) within a $3\arcsec$ radius. Entries flagged as “NOT GOOD” are excluded. After applying these procedures, the final training set consists of 8489 quasars and 5229 stars, as summarized in Table 2. The remaining 22,434 sources in the CSC–WISE–GAIA sample lack reliable spectroscopic classifications and are treated as unlabeled objects for subsequent prediction.
@@
QSO 61.88 8489
STAR 38.12 5229
Total training set 13,718
@@
Total 36,152
Although the training set is not strictly balanced, it provides broad coverage of photometric, astrometric, and X-ray parameter space for both quasars and stars. This construction is designed to reduce overfitting and to enable the RF classifier described in Section III to robustly distinguish quasars from stellar contaminants in the low-latitude sky.

### v2

@@ Abstract
Quasar samples remain severely incomplete at low Galactic latitudes because of strong extinction and source confusion. We conduct a systematic search for quasars behind the Galactic plane using X-ray sources from the Chandra Source Catalog (CSC 2.1), combined with optical data from Gaia DR3 and mid-infrared data from CatWISE2020. Using spectroscopically confirmed quasars and stellar-type objects from data sets including DESI, SDSS, and LAMOST, we apply a Random Forest classifier to identify quasar candidates, with stellar contaminants suppressed using Gaia proper-motion constraints. Photometric redshifts are estimated for the candidates using a Random Forest regression model. Applying this framework to previously unclassified CSC sources, we identify 7570 quasar candidates, including 1060 Galactic Plane Quasar (GPQ) candidates at $|b|<20^{\circ}$, of which 551 are high-confidence candidates. Relative to the previously known GPQ sample, our selected GPQs reach lower optical and X-ray fluxes, improving sensitivity to low-flux GPQs. In addition, both the GPQ candidates and known GPQs display harder X-ray spectra than the all-sky quasar sample, consistent with increased absorption through the Galactic plane. Pilot spectroscopy confirms two high-confidence GPQ candidates as quasars at spectroscopic redshifts of $z=1.2582$ and $z=1.1313$, and further spectroscopic follow-up of the GPQ sample is underway. This work substantially improves the census of GPQs and provides a valuable target sample for future spectroscopic follow-up, enabling the use of GPQs to refine the reference frames for astrometry and probe the Milky Way interstellar and circumgalactic media with the absorption features of GPQs.
@@ I INTRODUCTION
Quasars are among the most luminous active galactic nuclei (AGNs), powered by accretion onto supermassive black holes, and are detectable out to cosmological distances. Their high luminosities, compactness, and long-lived activity phases make quasars uniquely valuable probes for a wide range of astrophysical and cosmological questions, including the growth of supermassive black holes and their coevolution with host galaxies (e.g., Di Matteo et al. 2005; Kormendy and Ho 2013; Wu et al. 2015; Bañados et al. 2018; Inayoshi et al. 2020; Fan et al. 2023), and the large-scale structure of the Universe through quasar clustering and baryon acoustic oscillations (e.g., Eisenstein et al. 2005; Dawson et al. 2013; Blanton et al. 2017). In addition, quasars have negligible parallaxes and proper motions, making them ideal reference objects for defining and maintaining celestial reference frames (CRFs; e.g., Ma et al. 2009; Gaia Collaboration et al. 2022, 2023a). Despite major progress in wide-area surveys, however, the completeness of quasar samples remains poor close to the Galactic plane.
@@
Over the past decades, major spectroscopic surveys have dramatically expanded the known quasar population. The Sloan Digital Sky Survey (SDSS; Almeida et al., 2023; York et al., 2000) and its extensions established the first homogeneous, large-area quasar samples (Soria and Kong, 2002; Schneider et al., 2010; Pâris et al., 2017; Lyke et al., 2020), while complementary efforts such as the 2dF QSO Redshift Survey and the LAMOST quasar survey extended quasar identification to different depths and sky regions (Croom et al., 2004, 2009; Ai et al., 2016; Jin et al., 2023; Lyu et al., 2025). More recently, the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al., 2025) has delivered the largest spectroscopic quasar samples to date, with more than one million confirmed quasars now available for statistical studies. Despite this rapid progress, the sky distribution of known quasars remains highly non-uniform.
@@
In particular, the census of quasars at low Galactic latitudes remains severely incomplete. Most optical and near-infrared surveys avoid regions close to the Galactic plane, where strong interstellar extinction, high stellar densities, and severe source confusion substantially degrade the effectiveness of traditional color-based selection techniques (Schlegel et al., 1998; Ross et al., 2012; Richards et al., 2002). As a consequence, the surface density of spectroscopically confirmed quasars decreases sharply toward the Galactic plane, producing a persistent gap in all-sky quasar samples (Fu et al., 2021; Huo et al., 2025). This incompleteness limits the use of quasars as uniform tracers of large-scale structure and as astrometric reference objects, and motivates dedicated searches for quasars in the low-latitude sky.
@@
Quasars behind the Galactic plane (GPQs; $|b|\lesssim 20^{\circ}$), are nevertheless of considerable scientific importance. As distant extragalactic point sources with negligible parallaxes and proper motions, GPQs provide stable reference objects that are essential for constructing and validating celestial reference frames in the Galactic disk, where background sources are currently sparse (Gaia Collaboration et al. 2018b, a, 2023b). GPQs also serve as bright background beacons for absorption-line studies of the Milky Way interstellar and circumgalactic media, enabling measurements of gas kinematics, dust extinction, and chemical abundances along heavily obscured sightlines that are under-sampled at high Galactic latitudes (e.g., Savage and Sembach 1996; Richter 2006; Fu et al. 2021, 2024).
@@
Identifying GPQs is observationally challenging. The Galactic plane hosts a dense and diverse population of stars, binaries, and other Galactic sources, including many X-ray emitters, and these contaminants can overlap quasars in color–color and color–magnitude spaces. As a result, GPQs have been explored in relatively few dedicated studies. Early efforts by Im et al. (2007) uncovered dozens of bright QSOs/AGNs at $|b|<20^{\circ}$ using radio and near-infrared criteria. A major advance was achieved by Fu et al. (2021), who applied transfer learning to optical and infrared data from Pan-STARRS1 (PS1; Chambers et al., 2016) DR1 and AllWISE (Wright et al., 2010; Cutri et al., 2013), producing a catalog of roughly 160,000 GPQ candidates within $|b|<20^{\circ}$. Subsequent follow-up spectroscopy confirmed approximately 200 GPQs with high success rates (Fu et al., 2022). Werk et al. (2024) later confirmed 72 UV-bright AGNs at $|b|<30^{\circ}$, with 25 newly recognized objects, and the LAMOST GPQ program has recently reported $\sim 1300$ additional GPQs based on LAMOST DR10 (Huo et al., 2025). These studies demonstrate that combining multiwavelength data with modern classification approaches is essential for making progress in the Galactic plane.
@@
Machine-learning methods applied to multi-band photometry and astrometry can exploit high-dimensional information to improve quasar selection in crowded fields (Richards et al., 2002; Carrasco et al., 2015; Gaia Collaboration et al., 2023a; Fu et al., 2025). X-ray observations provide a powerful and complementary avenue for identifying quasars with extreme properties (Wang et al., 2025) or X-ray sources in the Galactic disk (Bao et al., 2025). Among current X-ray facilities, the Chandra X-ray Observatory is uniquely advantageous for low-latitude work: its sub-arcsecond angular resolution substantially mitigates source confusion in dense stellar regions, enabling more reliable associations between X-ray sources and their optical/infrared counterparts. The Chandra Source Catalog (CSC; here we use CSC 2.1) therefore offers an excellent basis for constructing GPQ candidates, provided that Galactic X-ray source contamination can be effectively controlled.
@@ II Data
We used Chandra Source Catalog version 2.1 (CSC 2.1; Evans et al., 2024) Master Sources Table to select GPQ candidates. It includes measured properties for 407,806 unique compact and extended X-ray sources and more than 1.3 million individual detections observed with either ACIS or HRC-I observations released publicly before the end of 2021. Most of these sources remain unidentified. Chandra has two focal plane instruments: Advanced CCD Imaging Spectrometer (ACIS) and High Resolution Camera (HRC). The ACIS instrument observes in broad (b): 0.5–7.0 keV, ultrasoft (u): 0.2–0.5 keV, soft (s): 0.5–1.2 keV, medium (m): 1.2–2.0 keV, and hard(h): 2.0–7.0 keV bands. The HRC instrument observes in 0.1–10 keV energy band and is designated as ‘W’ band. The energy flux in each band is determined using aperture photometry. The source count is derived from an elliptical source region and subtracted by the background count in the surrounding region. To convert the count rate to energy flux, the total count rate is summed up and then scaled by the local ancillary response function.
In this work, we use aperture-corrected average net-flux in b, u, s, m, and h bands, which are named as $b$-csc, $u$-csc, $s$-csc, $m$-csc, $h$-csc, respectively. We applied several constraints to the CSC 2.1 data to ensure quality. First, we selected only point sources with the extent flag set to 0 and a detection significance of likelihood >10>10. We then applied additional filters based on the following CSC 2.1 quality flags: pileup_flag, sat_src_flag, conf_flag, streak_src_flag (Table 1). Through this process, a total of 357,727 unique high-quality X-ray point sources were selected from CSC 2.1.
@@
We cross-match these 357,727 unique CSC 2.1 sources with the CatWISE2020 (Eisenhardt et al., 2020) and Gaia DR3 catalogues to obtain their multi-wavelength properties. Using the TOPCAT software (Taylor, 2005), we adopt matching radii of $3\arcsec$ and $1\farcs 5$ for CatWISE2020 and Gaia DR3, respectively, retaining only the nearest counterpart.
For the CatWISE2020 counterparts, we apply filters to retain sources with magnitudes in the ranges $8<W1<17.7$ and $7<W2<17.5$, and corresponding magnitude errors $\sigma\leq 0.2$, ensuring that the magnitudes are below the nominal limiting magnitudes in each band. For the Gaia DR3 counterparts, we require $G<21$ mag and a photometric error less than 0.2.
All Gaia magnitudes are then corrected for extinction using the two-dimensional dust reddening map from Planck Collaboration et al. (2016) and the Fitzpatrick (1999) extinction law assuming a total-to-selective extinction ratio of $R_{V}=3.1$. To mitigate the potential impact of bright sources on classification performance, we further excluded objects with extinction-corrected $G$-band magnitudes brighter than 10.
To improve the purity of the GPQ candidate sample, we then remove potential optically extended sources using the following procedure. We refine the point-source selection by applying the corrected BP/RP photometric excess factor, $C^{*}$, derived from Gaia DR3 (Riello et al., 2021), which provides an additional constraint on source morphology. The original BP/RP excess factor, $C$, is defined as the ratio of the summed BP and RP integrated fluxes to the $G$-band flux:
@@
Because the BP and RP photometric windows are broader than that of the $G$ band, extended sources tend to exhibit systematically larger excess factors than point sources (e.g., Liu et al., 2020). The corrected excess factor $C^{*}$ mitigates the color dependence inherent in $C$, providing a more reliable diagnostic of source extension (Riello et al., 2021). To compress the dynamic range and suppress the influence of extreme outliers, we impose the criterion $\log(1+C^{*})\leq 0.5$, which yields a cleaner point-source sample (e.g., Fu et al., 2025). Finally, 36,158 sources are selected with high-quality measurements from CSC 2.1, CatWISE2020, and Gaia DR3, among which 15,588 are located within the Galactic plane ($|b|<20^{\circ}$). Hereafter, we refer to this sample as the CSC–WISE–GAIA sample.
During the cross-matching procedure, a single X-ray source may be associated with multiple optical/infrared counterparts, leading to potential source confusion. This effect is particularly pronounced in the Galactic plane due to high extinction and stellar density. Indeed, among the 36,158 X-ray sources in our CSC–WISE–GAIA sample, 2,269 ($\sim$6%) possess multiple Gaia or CatWISE2020 counterparts within the matching radii. As expected, this confusion rate rises to $\sim$10% (1,527 out of 15,588 sources) for the subset located within the Galactic plane ($|b|<20^{\circ}$).
To evaluate the impact of source confusion, we conducted a parallel test using the probabilistic algorithm nway (Salvato et al., 2018). We found that while it improves sample purity, it does so at the expense of completeness (detailed in Section V.1). Because our primary objective is to retain a larger sample of quasar candidates in the Galactic plane, we adopted a nearest-neighbor matching strategy.
@@ II.2 The Training Set
A well-defined and representative training set is critical for the supervised classification of quasars and stellar-type objects, particularly in the crowded and heterogeneous environments near the Galactic plane. Following the construction of the CSC–WISE–GAIA sample described in Section II.1, we compile spectroscopically confirmed quasar and stellar samples from multiple surveys and catalogs to serve as labeled data for training the RF classifier.
Figure 1: Spatial distribution of the CSC-WISE-GAIA sample ($N=36,158$, pink) and the training set ($N=13,724$, blue) in Galactic coordinates using a Mollweide projection. The light green shaded region indicates the Galactic plane ($|b|<20^{\circ}$). This distribution demonstrates that while the training labels are primarily concentrated at high latitudes, our parent sample provides comprehensive coverage of the Galactic plane, enabling the discovery of new candidates in obscured regions.
We first identify quasars by cross-matching the CSC–WISE–GAIA sample with spectroscopically confirmed quasars from DESI DR1 (DESI Collaboration et al., 2025) and SDSS DR18 (Almeida et al., 2023), adopting a matching radius of $1\arcsec$. To improve completeness, this initial quasar sample is further supplemented with confirmed quasars from the Million Quasars Catalog (Flesch, 2023), the LAMOST spectroscopic surveys (Cui et al., 2012; Luo et al., 2015), and the Galactic Plane quasar sample identified from LAMOST DR10 by Huo et al. (2025). All supplementary quasar catalogs are cross-matched with the CSC–WISE–GAIA sample using a $3\arcsec$ radius to ensure consistent source associations.
For the stellar class, we similarly cross-match the CSC–WISE–GAIA sample with DESI DR1 and SDSS DR18 to identify spectroscopically classified stars. To better represent the diverse Galactic X-ray source population that may contaminate quasar selection, we further include several specialized stellar catalogs. These include Galactic Wolf–Rayet stars (van der Hucht, 2001, 2006) and APOGEE-2 stars from SDSS DR16 (Jönsson et al., 2020); young stellar objects (YSOs) in open clusters and molecular clouds (e.g., Ozawa et al., 2005; Giardino et al., 2007; Delgado et al., 2011; Povich et al., 2011; Rebull et al., 2011; Megeath et al., 2012); high-mass X-ray binaries (HMXBs) and related Galactic and nearby extragalactic sources (Liu et al., 2006; Walter et al., 2015; Mineo et al., 2012; Sazonov and Khabibullin, 2017); low-mass X-ray binaries (LMXBs) compiled from the Ritter–Kolb catalog and subsequent updates including recent comprehensive catalogs (Ritter and Kolb, 2015; Liu et al., 2007; Sazonov et al., 2020; Kundu et al., 2007; Humphrey and Buote, 2008; Zhang et al., 2011; Avakyan et al., 2023; Neumann et al., 2023); and cataclysmic variables (CVs) (Downes et al., 2005; Ritter and Kolb, 2015). These catalogs are cross-matched with the CSC–WISE–GAIA sample using a matching radius of $3\arcsec$. All sources matched to DESI/SDSS stars or identified in the aforementioned stellar catalogs are classified as stellar-type objects in this paper.
@@
QSO 61.86 8489
STAR 38.14 5235
Total Training Set 13,724
@@
Total No. 36,158
Note. “QSO” and “STAR” represent quasars and stellar-type objects, respectively.
Sources that are not matched to any of the above quasar or stellar catalogs are subsequently queried against the SIMBAD database (Wenger et al., 2000) within a $3\arcsec$ radius. Entries flagged as “NOT GOOD” are excluded. After applying these procedures, the final training set consists of 8489 quasars and 5235 stellar-type objects, as summarized in Table 2. The remaining 22,434 sources in the CSC–WISE–GAIA sample lack reliable spectroscopic classifications and are treated as unlabeled objects for subsequent prediction.
The spatial distribution of the final datasets is illustrated in Figure 1. We present the $N=36,158$ sources in the parent CSC-WISE-GAIA sample (pink) alongside the $N=13,724$ spectroscopically confirmed sources in the training set (blue) using a Mollweide projection in Galactic coordinates. While the training set is predominantly concentrated at high Galactic latitudes due to the footprint of major spectroscopic surveys like SDSS and DESI, our parent sample exhibits comprehensive coverage across the entire sky, including the Galactic plane ($|b|<20^{\circ}$, shaded in light green). This visualization highlights that our machine-learning approach is strategically positioned to identify new candidates in regions where spectroscopic labels are currently sparse.
Although the training set is not strictly balanced, it provides broad coverage of photometric, astrometric, and X-ray parameter space for both quasars and stellar-type objects. This construction is designed to reduce overfitting and to enable the RF classifier described in Section III to robustly distinguish quasars from stellar contaminants in the low-latitude sky.